OpenStudy (anonymous):

Solve the system using any method (graphing, elimination, or substitution): (1) 2x^2 + y^2 = 18 (2) xy = 4 I know I can't use elimination, and I don't know how I would graph the second one, so I'm trying substitution. Am I doing it right to far? xy = 4 y = 4/x then substitute into equation (1) 2x^2 + (4/x)^2 = 18 2x^2 + 16/(x^2) = 18 If I'm right, what's the next step? I'm a bit confused.

OpenStudy (anonymous):

From where you stopped.. => \[2x^{2}+ 16/x^2 = 18\] \[2x^4-18x^2+16 = 0\] \[x^4-9x^2+8=0\] Should be able to solve from there!

OpenStudy (anonymous):

Thank you!